OpenStudy (anonymous):

differentiate the function f(t)=4+ln(t)/6-ln(t)

OpenStudy (tkhunny):

AS WRITTEN, namely, \(f(t) = 4 + \dfrac{ln(t)}{6} - ln(t)\), we have \(f'(t) = \dfrac{1}{6t} - \dfrac{1}{t} = \dfrac{-5}{6t}\)... HOWEVER, I suspect you intended to write \(f(t) = \dfrac{4+\ln(t)}{6+\ln(t)}\). If so, how's your Quotient Rule working?

OpenStudy (anonymous):

The quotient rule our teacher taught is (low derivativehigh-high dlow)/(the square of whats below) if that makes any sense